OpenStudy (anonymous):

Is there a simple way to copy, paste, and edit an equation someone else has posted? This would be quite desirable.

OpenStudy (kinggeorge):

Right click on the equation, go to "Show Math As" and click on "TeX Commands"

OpenStudy (anonymous):

Hmm haven't I tried that before? Testing now... \[\frac{\sqrt{\sin{x^2}}}{\cos{x^3}^5}\]

OpenStudy (anonymous):

So frustrating. It opens up a new chrome window for me, but that window is not viewable...

OpenStudy (amistre64):

yeah, it used to pop up a new window with the eq in it; now it just hides it and makes you go searching for it ...

OpenStudy (farmdawgnation):

Yeah, there's a bug in Chrome. There's not much we can do about it until the MathJAX guys fix that (it's their code not ours). :(

OpenStudy (kinggeorge):

The bug doesn't appear in all chrome versions. It works fine for the beta version.

OpenStudy (amistre64):

its almost as if its opening it and positioning it in a negative screen space.

OpenStudy (anonymous):

Right. The workaround that I figured out, in case anyone is interested, is to right click the window and select "Show as tab."

OpenStudy (farmdawgnation):

@KingGeorge - Do you mean Chrome Canary? If so, that's great! I should test that

OpenStudy (kinggeorge):

Nope just the most recent beta 19.0.1084.41 beta-m

OpenStudy (farmdawgnation):

Epic, that means that it'll be resolved in the next release. Good news! :)